Please solve by addition method:
3x - 4y = 9
-3x + 4y = 12

Answer by rapaljer(4671)

If you just add these two equations together, you will subtract out the x terms AND the y terms:
3x - 4y = 9
-3x + 4y = 12

0 + 0 = 21
Obviously this is NOT possible. There is NO way that 0 = 21! At least not in OUR lifetime!! This means that there is NO SOLUTION. Some textbooks will call this the Empty Set, and they write this either with empty set braces: { }, or with the Greek letter "phi" which is a 0 with a slash: +phi+. I personally think this is very confusing to students, so I prefer to just call it NO SOLUTION.

Now, think about what situtation could happen with equations of two straight lines that would cause a NO SOLUTION situation. What happened is that the lines are PARALLEL!!

R^2 at SCC